# Upper bound for the total mean curvature of spin fill-ins

### v2

@@ Abstract.
Gromov conjectured that the total mean curvature of the boundary of a compact Riemannian manifold can be estimated from above by a constant depending only on the boundary metric and on a lower bound for the scalar curvature of the fill-in. We prove Gromov’s conjecture if the manifolds are spin and the mean curvature $H$ is non-negative. One can also allow $H$ to take negative values, but then the constant depends on a (negative) lower bound for $H$.
@@ 1. Introduction
Of course, $M$ may not bound any compact manifold $X$, but if it does so, it follows from [SWW]*Theorem 1.1 by Shi, Wang, and Wei that the given metric on $M$ can be extended to a metric on $X$ with $\mathrm{scal}_{X}>0$. Their result gives no control on the mean curvature of the boundary. Theorem 3.7 in [BH] by the author and Hanke implies that the metric on $X$ can be deformed in such a way that its scalar curvature remains positive, it still induces the given metric on $M$, and the mean curvature of the boundary becomes arbitrarily negative. Thus
@@
Gromov conjectured that $\int_{M}H$ cannot be made arbitrarily large. More precisely, the following is believed to hold:
@@
We prove Gromov’s conjecture for spin manifolds if the mean curvature is non-negative. More precisely, we show the following:
Theorem.
@@
$\int_{M}H\leq C(M)+\Big(\sqrt{\tfrac{n}{n+1}}\,\lambda+\eta\Big)\mathsf{vol}(M),$
@@
All previously known results need the assumption that $H\geq 0$ or even $H>0$. If $\lambda=0$, $H>0$, $X$ is spin, and each connected component of $M$ admits an isometric embedding as a strictly convex hypersurface in $\mathbb{R}^{n+1}$, then this has been proved by Shi and Tam in [ST]*Theorems 1 and 4.1. The constant $C(M)$ is in this case given by the total mean curvature of the isometric embedding of $M$ into $\mathbb{R}^{n+1}$. The condition that each connected component of $M$ embeds isometrically as a strictly convex hypersurface into $\mathbb{R}^{n+1}$ has been relaxed by Eichmair, Miao, and Wang in [EMW] to require the components to have non-negative scalar curvature and admit embeddings as star-shaped hypersurfaces.
The case where $M$ is diffeomorphic to a sphere with $\lambda=0$, $H>0$, and $X$ spin has been treated by Shi, Wang, and Wei in [SWW]*Theorem 4.1.
Mantoulidis and Miao investigated in [MM] the case where $n=2$ and $M$ has arbitraty topology, again if $\lambda=0$ and $H>0$.
If $M=T^{n}$ is a torus, $X=D^{2}\times T^{n-1}$, $n\leq 6$, and $H>0$, then Gromov’s conjecture has been proved by Wang in [W]*Theorems 1.6 and 1.8. In this case, $\lambda$ is allowed to be negative. If the torus $M$ carries a flat metric, then $C(M)$ is related to the systole of $M$.
@@
In the mean time, Frenck, Hanke, and Hirsch have presented a different proof of Gromov’s conjecture if $H\geq 0$, using surgery and deformation methods ([FHH]). Their approach has the advantage that it also works in the non-spin case if $n\leq 6$.
@@ 2. A boundary value problem
Let $X$ be a compact connected Riemannian spin manifold of dimension $n+1\geq 2$ with boundary $\partial X=M$. The spinor bundle $\Sigma X$ of $X$ is a Hermitian vector bundle with metric connection. We choose the convention that the fiberwise scalar product of $\Sigma X$ is antilinear in the first argument and linear in the second argument. Let $\gamma\colon TX\to\operatorname{\mathsf{End}}(\Sigma X)$ denote the Clifford multiplication. It satisfies the Clifford relations
@@
We denote the space of square integrable sections of a vector bundle $E$ over $X$ by $L^{2}(X,E)$ and the $L^{2}$-Sobolev space of order $\alpha$ by $H^{\alpha}(X,E)$ and similarly for $M$.
@@ 3. Some preparation
It is well known that there exists a smooth partition of unity $\psi_{j}\in C^{\infty}(M,\mathbb{R})$ subordinate to the open cover, i.e., $\psi_{j}\geq 0$ and $\mathrm{supp}(\psi_{j})\subset U_{j}$ for every $j$ and $\sum_{j}\psi_{j}\equiv 1$ on $M$. Then
@@
4. Proof of the theorem
In this section we carry out the proof of the theorem stated in the introduction.
@@ 4.1. Definition of the constant C(M)
$C(M):=4\sum_{j=1}^{m}\|D_{M}\varphi_{j}\|_{L^{2}(M)}\cdot\|\varphi_{j}\|_{L^{2}(M)}.$
@@ 4.2. The estimate
Now let $X$ be a spin fill-in of $M$, i.e., $M=\partial X$ as a Riemannian spin manifold. Without loss of generality, we may assume that $X$ is connected since otherwise we can apply the following argument to each connected component separately. Connected components of $X$ with empty boundary do not contribute to the integral of the mean curvature and can thus be ignored.
@@
$\nabla_{\nu}\Phi_{j}$ $=-\gamma(\nu)D\Phi_{j}-D_{M}\Phi_{j}+\tfrac{1}{2}H\Phi_{j}=i\,s\,D\Phi_{j}-D_{M}\Phi_{j}+\tfrac{1}{2}H\Phi_{j}=-\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\,s\,\Phi_{j}-D_{M}\Phi_{j}+\tfrac{1}{2}H\Phi_{j}.$
@@
$\int_{M}H\leq C(M)+\bigg(\sqrt{\frac{n}{n+1}}\lambda+\eta\bigg)\mathsf{vol}(M).$
This completes the proof of the theorem.
@@ 5. Concluding remarks
Example.
For $\kappa\in\mathbb{R}$ let $\mathbb{M}_{\kappa}^{n+1}$ be the $(n+1)$-dimensional simply connected space form of constant sectional curvature $\kappa$. Let $X\subset\mathbb{M}_{\kappa}^{n+1}$ be a compact ball where the radius is chosen such that $M=\partial X$ is the standard unit sphere. This is possible if $\kappa\leq 1$. The mean curvature of $M$ is given by $H=n\sqrt{1-\kappa}$. For negative $\kappa$ we set $n(n+1)\kappa=-\lambda^{2}$ with $\lambda>0$. Then $\mathrm{scal}_{X}=-\lambda^{2}$ and we have
$\int_{M}H=n\cdot\sqrt{1+\tfrac{\lambda^{2}}{n(n+1)}}\cdot\mathsf{vol}(M)=\sqrt{\tfrac{n^{2}}{\lambda^{2}}+\tfrac{n}{n+1}}\cdot\lambda\cdot\mathsf{vol}(M).$
This shows that the dependence on $\lambda$ in our estimate is of the right order as $\lambda\to\infty$.
@@
The definition of the constant $C(M)$ in Section 4.1 depends on the spin structure of $M$. However $M$, being compact, has only finitely many spin structures. Thus, $C(M)$ can be chosen independently of the spin structure.

### v3

@@ Abstract.
Gromov conjectured that the total mean curvature of the boundary of a compact Riemannian manifold can be estimated from above by a constant depending only on the boundary metric and on a lower bound for the scalar curvature of the fill-in. We prove Gromov’s conjecture if the manifolds are spin with a constant that also depends on a lower bound on the mean curvature $H$ (which is allowed to take negative values). If the boundary is a (not necessarily convex) hypersurface in a space form of non-negative curvature, then the constant can be made explicit in terms of the mean curvature of this model embedding. If the boundary has constant sectional curvature $\kappa>0$ and is a projective space of dimension $n\equiv 3\mod 4$ or a sphere, then the constant can be expressed in terms of $\kappa$. If the boundary is a flat torus, then the constant can be expressed in terms of lattice data.
@@ 1. Introduction
Of course, $M$ may not bound any compact manifold $X$, but if it does, it follows from [SWW]*Theorem 1.1 by Shi, Wang, and Wei that the given metric on $M$ can be extended to a metric on $X$ with $\mathrm{scal}_{X}>0$. Their result gives no control over the boundary mean curvature. Theorem 3.7 in [BH] by the author and Hanke implies that the metric on $X$ can be deformed in such a way that its scalar curvature remains positive, it still induces the given metric on $M$, and the mean curvature of the boundary becomes arbitrarily negative. Thus
@@
While we can make $H$ as small as we like, Gromov conjectured that $\int_{M}H$ cannot be made arbitrarily large. More precisely, the following is believed to hold:
@@
We prove Gromov’s conjecture for spin manifolds with a constant that also depends on a lower bound for the mean curvature. More precisely, we show the following:
Theorem 1.
@@
$\frac{1}{\mathsf{vol}(M)}\int_{M}H\leq C(M)+\sqrt{\tfrac{n}{n+1}}\,\lambda+\eta,$
@@
In Example 2, we choose $X$ to be a ball in a space form and see that the term $\sqrt{\tfrac{n}{n+1}}\,\lambda$ gives the precise asymptotics for $\lambda\to\infty$. Thus, the coefficient $\sqrt{\tfrac{n}{n+1}}$ is optimal.
Most previously known results require the assumption that $H\geq 0$ or even $H>0$. The case where $M$ is diffeomorphic to a sphere with $\lambda=0$, $H>0$, and $X$ spin has been treated by Shi, Wang, and Wei in [SWW]*Theorem 4.1. Mantoulidis and Miao investigated in [MM] the case where $n=2$ and $M$ has arbitrary topology, again if $\lambda=0$ and $H>0$. In these results, the constant $C(M,\lambda)$ is not explicit.
@@
Using entirely different surgery and deformation methods, Frenck, Hanke, and Hirsch have presented an independent proof of Gromov’s conjecture with lower mean curvature bound in [FHH]. Their estimate is of the form
$\frac{1}{\mathsf{vol}(M)}\int_{M}H\leq C(M)(1+\lambda+\eta)$
which also gives linear growth in $\lambda$ and $\eta$, but with unknown coefficients. Their approach has the advantage that it allows one to drop the spin assumption if $n\leq 6$ and $\eta=0$. The dimension restriction is due to the use of the hyperbolic positive mass theorem.
If we make additional assumptions on $M$, the constant $C(M)$ can be made explicit. For $\kappa\in\mathbb{R}$, let $\mathbb{M}_{\kappa}^{n+1}$ be the $(n+1)$-dimensional simply connected space form of constant sectional curvature $\kappa$. For $\kappa=0$, this is the Euclidean space $\mathbb{R}^{n+1}$, for $\kappa=1$ the round unit sphere $S^{n+1}$, and for $\kappa=-1$ the hyperbolic space $\mathbb{H}^{n+1}$.
Theorem 2.
In addition to the assumptions in Theorem 1, assume that $M$ is isometric to an oriented hypersurface of $\mathbb{M}_{\kappa}^{n+1}$ where $\kappa\geq 0$. Denote the (unnormalized) mean curvature of this embedding by $H_{0}$. Then we have:
$\frac{1}{\mathsf{vol}(M)}\int_{M}H\leq 2\,\sqrt{\frac{1}{\mathsf{vol}(M)}\int_{M}H_{0}^{2}}+2n\sqrt{\kappa}+\sqrt{\tfrac{n}{n+1}}\,\lambda+\eta.$
Example 1.
The case $\kappa=\lambda=0$ and $H>0$ has been studied by Shi and Tam in [ST]*Theorem 4. Our estimate states that
$\frac{1}{\mathsf{vol}(M)}\int_{M}H\leq 2\,\sqrt{\frac{1}{\mathsf{vol}(M)}\int_{M}H_{0}^{2}}.$ (2)
while Shi and Tam proved the sharper estimate
$\int_{M}H\leq\int_{M}H_{0}$
under the additional assumption that each connected component of $M$ is a strictly convex hypersurface in $\mathbb{R}^{n+1}$. This strict convexity has been relaxed by Eichmair, Miao, and Wang in [EMW] to require the components to have non-negative scalar curvature and admit embeddings as star-shaped hypersurfaces.
We do not need any convexity assumption on $M$. In particular, the components of $M$ need not be diffeomorphic to a sphere. Moreover, we only need $H\geq 0$ rather than $H>0$ to obtain (2).
Theorem 3.
In addition to the assumptions in Theorem 1, assume that $M$ is isometric to one of the following:
  • a flat torus,
  • a round sphere,
  • round projective space $\mathbb{R}P^{n}$ with $n\equiv 3\mod 4$.
Denote the eigenvalue of the Dirac operator of $M$ of smallest modulus by $\mu_{1}$. Then we have:
$\frac{1}{\mathsf{vol}(M)}\int_{M}H\leq 4\,|\mu_{1}|+\sqrt{\tfrac{n}{n+1}}\,\lambda+\eta.$
The Dirac eigenvalue $\mu_{1}$ is explicitly known in the cases mentioned above. If $M$ is a round sphere or a round projective space of sectional curvature $\kappa$, then $|\mu_{1}|=\frac{n}{2}\sqrt{\kappa}$. If $M$ is a flat torus, then $|\mu_{1}|$ can be expressed in terms of lattice data; see [Fr]*Satz 2 for details. The $n$-dimensional torus has $2^{n}$ spin structures, and for one of them we have that $\mu_{1}=0$. This is precisely the spin structure for which the torus does not bound as a spin manifold.
Wang has proved an estimate for flat tori in terms of a spin systole in [W]*Theorem 1.6 under the assumption that $\lambda=\sqrt{n(n+1)}$, $H>0$, $2\leq n\leq 6$, and $X$ is diffeomorphic to a $2$-disk times a torus.
The paper is structured as follows: In Section 2 we prove the well-posedness of a boundary value problem for the Dirac operator. This may be of independent interest. Section 3 contains some technical preparation. The next three sections contain the proofs of the theorems. The main argument is contained in Section 4. The proofs of Theorems 2 and 3 improve the constant $C(M)$ in the special situation under consideration and make it explicit. The final section contains the example of balls in a space form and some concluding remarks.
@@ 2. A boundary value problem
Let $X$ be a compact connected Riemannian spin manifold of dimension $n+1\geq 2$ with boundary $\partial X=M$. Let $M$ carry the metric and the spin structure induced from $X$. The spinor bundle $\Sigma X$ of $X$ is a Hermitian vector bundle with metric connection. We choose the convention that the fiberwise scalar product of $\Sigma X$ is antilinear in the first argument and linear in the second argument. Let $\gamma\colon TX\to\operatorname{\mathsf{End}}(\Sigma X)$ denote the Clifford multiplication. It satisfies the Clifford relations
@@
We denote by $L^{2}(X,E)$ the space of square-integrable sections of a vector bundle $E$ over $X$ and by $H^{\alpha}(X,E)$ the $L^{2}$-Sobolev space of order $\alpha$, and similarly for $M$.
@@ 3. Some preparation
It is well known that there is a smooth partition of unity $\psi_{j}\in C^{\infty}(M,\mathbb{R})$ subordinate to the open cover, i.e., $\psi_{j}\geq 0$ and $\mathrm{supp}(\psi_{j})\subset U_{j}$ for every $j$ and $\sum_{j}\psi_{j}\equiv 1$ on $M$. Then
@@
4. Proof of Theorem 1
In this section we carry out the proof of Theorem 1.
@@ 4.1. Definition of the constant C(M)
$C(M):=\frac{4}{\mathsf{vol}(M)}\cdot\sum_{j=1}^{m}\|D_{M}\varphi_{j}\|_{L^{2}(M)}\cdot\|\varphi_{j}\|_{L^{2}(M)}.$ (9)
@@ 4.2. The estimate
Now let $X$ be a spin fill-in of $M$, i.e., $M=\partial X$ as a Riemannian spin manifold. Without loss of generality, we may assume that $X$ is connected, as otherwise we can apply the following argument to each connected component separately. Connected components of $X$ with empty boundary do not contribute to the integral of the mean curvature and can thus be ignored.
@@
$\nabla_{\nu}\Phi_{j}$ $=-\gamma(\nu)D\Phi_{j}-D_{M}\Phi_{j}+\tfrac{1}{2}H\Phi_{j}$
$=i\,s\,D\Phi_{j}-D_{M}\Phi_{j}+\tfrac{1}{2}H\Phi_{j}$
$=-\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\,s\,\Phi_{j}-D_{M}\Phi_{j}+\tfrac{1}{2}H\Phi_{j}.$
@@
$\int_{M}H\leq\bigg(C(M)+\sqrt{\frac{n}{n+1}}\lambda+\eta\bigg)\mathsf{vol}(M).$
This completes the proof of Theorem 1.
5. Proof of Theorem 2
In order to obtain Theorem 2, we need to refine the proof of Theorem 1, improving the constant $C(M)$ and making it explicit. Let $h\in C^{\infty}(M,\mathbb{R})$, let $\tau$ be a smooth field of pointwise isometries of $E$, and let $\beta\geq 0$ be a constant.
Lemma 4.
Assume the situation of the proof of Theorem 1. If $\varphi_{1},...,\varphi_{m}$ can be chosen such that for each $j$
$D_{M}\varphi_{j}=\varepsilon_{j}h\varphi_{j}+\beta\tau\varphi_{j}.$
holds, where $\varepsilon_{j}\in\{-1,1\}$, then we have
$\sum_{j=1}^{m}\int_{M}\langle D_{M}\,\Phi_{j},\Phi_{j}\rangle\leq 2\,\mathsf{vol}(M)\Bigg(\sqrt{\frac{1}{\mathsf{vol}(M)}\int_{M}h^{2}}+\beta\Bigg).$
Proof.
As in the proof of Theorem 1, $\varphi_{1},...,\varphi_{m}$ are scaled such that (8) holds. Define the orthogonal projection $P_{\tau}:=\tau^{-1}P^{-}\tau$. We compute, for any constants $\alpha_{1},\alpha_{2}>0$:
$\sum_{j=1}^{m}\int_{M}\langle D_{M}\,\Phi_{j},\Phi_{j}\rangle$ $=2\mathsf{Re}\sum_{j=1}^{m}\int_{M}\langle D_{M}\,\varphi_{j},P^{-}\Phi_{j}\rangle$
$=2\mathsf{Re}\sum_{j=1}^{m}\int_{M}\langle\varepsilon_{j}h\varphi_{j}+\beta\tau\varphi_{j},P^{-}\Phi_{j}\rangle$
$=2\mathsf{Re}\sum_{j=1}^{m}\int_{M}\langle\varepsilon_{j}hP^{-}\varphi_{j}+\beta P^{-}\tau\varphi_{j},P^{-}\Phi_{j}\rangle$
$=2\mathsf{Re}\sum_{j=1}^{m}\int_{M}\langle\varepsilon_{j}hP^{-}\varphi_{j}+\beta\tau P_{\tau}\varphi_{j},P^{-}\Phi_{j}\rangle$
$\leq 2\sum_{j=1}^{m}\big(\|hP^{-}\varphi_{j}\|_{L^{2}(M)}+\beta\|\tau P_{\tau}\varphi_{j}\|_{L^{2}(M)}\big)\|P^{-}\Phi_{j}\|_{L^{2}(M)}$
$=2\sum_{j=1}^{m}\big(\|hP^{-}\varphi_{j}\|_{L^{2}(M)}+\beta\|P_{\tau}\varphi_{j}\|_{L^{2}(M)}\big)\|P^{-}\Phi_{j}\|_{L^{2}(M)}$
$\leq 2\sum_{j=1}^{m}\big(\|hP^{-}\varphi_{j}\|_{L^{2}(M)}+\beta\|P_{\tau}\varphi_{j}\|_{L^{2}(M)}\big)\|P^{+}\Phi_{j}\|_{L^{2}(M)}$
$=2\sum_{j=1}^{m}\big(\|hP^{-}\varphi_{j}\|_{L^{2}(M)}+\beta\|P_{\tau}\varphi_{j}\|_{L^{2}(M)}\big)\|P^{+}\varphi_{j}\|_{L^{2}(M)}$
$\leq\sum_{j=1}^{m}\big(\alpha_{1}\|hP^{-}\varphi_{j}\|^{2}+\tfrac{1}{\alpha_{1}}\|P^{+}\varphi_{j}\|^{2}+\beta^{2}\alpha_{2}\|P_{\tau}\varphi_{j}\|^{2}+\tfrac{1}{\alpha_{2}}\|P^{+}\varphi_{j}\|^{2}\big)$
$=\alpha_{1}\int_{M}h^{2}+(\tfrac{1}{\alpha_{1}}+\beta^{2}\alpha_{2}+\tfrac{1}{\alpha_{2}})\mathsf{vol}(M).$
Choosing $\alpha_{1}=\sqrt{\mathsf{vol}(M)/\int_{M}h^{2}}$ and $\alpha_{2}=1/\beta$ yields [^2] If $h$ vanishes identically, then consider the limit $\alpha_{1}\to\infty$.
$\sum_{j=1}^{m}\int_{M}\langle D_{M}\,\Phi_{j},\Phi_{j}\rangle\leq 2\sqrt{\mathsf{vol}(M)\int_{M}h^{2}}+2\beta\mathsf{vol}(M).\qed$
Now we prove Theorem 2. For $\kappa\geq 0$, the spinor bundle of $\mathbb{M}_{\kappa}^{n+1}$ can be trivialized by Killing spinors, i.e., spinors $\Psi_{1},...,\Psi_{m}$ satisfying
$\nabla_{X}\Psi_{j}=\frac{\sqrt{\kappa}}{2}\gamma(X)\Psi_{j}.$
These Killing spinors are pointwise orthonormal if chosen appropriately. [^3] For $\kappa<0$, the spinor bundle of $\mathbb{M}_{\kappa}^{n+1}$ can also be trivialized by Killing spinors (with imaginary constant) but they are not pointwise orthonormal.
Now let $M\subset\mathbb{M}_{\kappa}^{n+1}$ be an oriented hypersurface. We restrict the Killing spinors $\Psi_{j}$ to $M$ and obtain a global frame $\varphi_{j}:=\Psi_{j}|_{M}$ of the spinor bundle $E=\Sigma\mathbb{M}_{\kappa}^{n+1}|_{M}\to M$. We rescale this frame as in Section 4.1 so that (8) holds. We choose one unit normal field $\nu_{0}$ for $M$ in $\mathbb{M}_{\kappa}^{n+1}$ and denote the corresponding mean curvature by $H_{0}$. Using (4), we compute
$D_{M}\varphi_{j}$ $=\tfrac{1}{2}H_{0}\varphi_{j}-\gamma(\nu_{0})D\Psi_{j}-\nabla_{{\nu_{0}}}\Psi_{j}$
$=\tfrac{1}{2}H_{0}\varphi_{j}+(n+1)\frac{\sqrt{\kappa}}{2}\gamma(\nu_{0})\Psi_{j}-\frac{\sqrt{\kappa}}{2}\gamma(\nu_{0})\Psi_{j}$
$=\tfrac{1}{2}H_{0}\varphi_{j}+\frac{n\sqrt{\kappa}}{2}\gamma(\nu_{0})\varphi_{j}.$
We apply Lemma 4 with $h=\frac{H_{0}}{2}$, $\varepsilon_{1}=\cdots=\varepsilon_{m}=1$, $\beta=\frac{n\sqrt{\kappa}}{2}$, and $\tau=\gamma(\nu_{0})$ to obtain
$\sum_{j=1}^{m}\int_{M}\langle D_{M}\,\Phi_{j},\Phi_{j}\rangle\leq\mathsf{vol}(M)\Bigg(\sqrt{\frac{1}{\mathsf{vol}(M)}\int_{M}H_{0}^{2}}+n\sqrt{\kappa}\Bigg)=\tfrac{1}{2}\,C^{\prime}(M)\,\mathsf{vol}(M),$
where we have put $C^{\prime}(M)=2\sqrt{\frac{1}{\mathsf{vol}(M)}\int_{M}H_{0}^{2}}+2n\sqrt{\kappa}$. Now the proof of Theorem 1 carries over with $C(M)$ replaced by $C^{\prime}(M)$. This yields Theorem 2.
6. Proof of Theorem 3
For all manifolds listed in the statement of Theorem 3, the eigenspace of $\tilde{D}_{M}^{2}$ for the smallest eigenvalue $\mu_{1}^{2}$ has a basis consisting of pointwise orthonormal eigenspinors, see [Fr] for the case of flat tori and [Baer1]*Section 4 for the spherical cases. This eigenspace is the orthogonal sum of the eigenspaces of $\tilde{D}_{M}$ itself for the eigenvalues $\mu_{1}$ and $\mu_{1}$. Thus, the pointwise orthonormal basis can be chosen to consist of eigenspinors of $\tilde{D}_{M}$. This yields a pointwise orthonormal basis $\varphi_{1},...,\varphi_{m}$ of $E$, consisting of eigensections of $D_{M}$ for the eigenvalues $\pm\mu_{1}$ by doubling if $n$ is odd.
We rescale this basis as in Section 4.1 so that (8) holds. We apply Lemma 4 with $h=\mu_{1}$ and $\beta=0$ to obtain
$\sum_{j=1}^{m}\int_{M}\langle D_{M}\,\Phi_{j},\Phi_{j}\rangle\leq 2\,\mathsf{vol}(M)\sqrt{\frac{1}{\mathsf{vol}(M)}\int_{M}\mu_{1}^{2}}=2\,|\mu_{1}|\,\mathsf{vol}(M)=\tfrac{1}{2}C^{\prime\prime}(M)\,\mathsf{vol}(M),$
where $C^{\prime\prime}(M)=4|\mu_{1}|$. Now the proof of Theorem 1 carries over with $C(M)$ replaced by $C^{\prime\prime}(M)$. This yields Theorem 3.
@@ 7. Concluding remarks
Example 2.
Let $X\subset\mathbb{M}_{\kappa}^{n+1}$ be a compact ball where the radius is chosen such that $M=\partial X$ is the standard unit sphere. This is possible if $\kappa\leq 1$. The mean curvature of $M$ is given by $H=n\sqrt{1-\kappa}$. For negative $\kappa$ we set $n(n+1)\kappa=-\lambda^{2}$ with $\lambda>0$. Then $\mathrm{scal}_{X}=-\lambda^{2}$ and we have
$\frac{1}{\mathsf{vol}(M)}\int_{M}H=n\cdot\sqrt{1+\tfrac{\lambda^{2}}{n(n+1)}}=\sqrt{\tfrac{n^{2}}{\lambda^{2}}+\tfrac{n}{n+1}}\cdot\lambda\sim\sqrt{\tfrac{n}{n+1}}\cdot\lambda\quad\text{ as }\lambda\to\infty.$
This shows that the dependence on $\lambda$ in our estimate is of the right order for large $\lambda$.
@@
The definition of the constant $C(M)$ in Section 4.1 depends on the spin structure of $M$. However, since $M$ is compact, it has only finitely many spin structures. Thus, $C(M)$ can be chosen independently of the spin structure.